# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 4. Polynomial Property and the Monodromy Groups
$\mathrm{Stab}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}.$ (4.2)
To this end, we first need to prove Theorem 4.2 which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
4.1. Periods of the generalized Frobenius manifold
Let us consider flat coordinates of the flat metric $g+\lambda\eta$ of $\overline{\mathcal{M}}$. They satisfy the equations
$\nabla_{g+\lambda\eta}\mathrm{d}u(\lambda)=0$ (4.3)
@@ 4.1. Periods of the generalized Frobenius manifold
$M_{\lambda}=(\overline{\mathcal{M}}\times\mathbb{C}^{\times})\setminus{\Sigma_{\lambda}},$ (4.4)
@@
${\Sigma}_{\lambda}=\{(z,\lambda)\in M\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
@@
$\mathscr{Q}(\widehat{M_{\lambda}})=\mathscr{O}(\widehat{M_{\lambda}})/\mathbb{C}.$
In the coordinates $(z^{1},\dots,z^{\ell})$, the equation (4.3) reads
$\frac{\partial^{2}u}{\partial z^{i}\partial z^{j}}-\Gamma_{\lambda,ij}^{k}\frac{\partial u}{\partial z^{k}}=0,$ (4.5)
Here $\Gamma_{\lambda,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$. Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$. By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ such that
$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}\right).$ (4.6)
@@
depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are homogeneous of degree zero, i.e., they satisfy the equation
$\kappa\lambda\frac{\partial u}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial u}{\partial z^{j}}=0.$ (4.8)
We denote by $\mathcal{S}_{\lambda}$ the space of periods of $\overline{\mathcal{M}}$ which are homogeneous of degree zero, then $\dim\mathcal{S}_{\lambda}=\ell$, and
$\mathcal{S}_{\lambda}=\mathrm{span}_{\mathbb{C}}\left\{\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)\right\}.$ (4.9)
Now we proceed to find another basis of $\mathcal{S}_{\lambda}$ that is associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
Proposition 4.1.
For each homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a solution
$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ (4.10)
of the equations (4.5) and (4.8), where $h^{\alpha}(z,\rho)$ can be analytically continued to $\rho=0$ such that $h^{\alpha}(z,0)=t^{\alpha}(z)$.
@@
Recall that we made the assumption in Remark 3.1 that $d_{\alpha}>0$, so from (3.20) we have $0<d_{\alpha}<\kappa$. Let us looking for a solution of (4.5) and (4.8) of the form $u(z,\rho)=\rho^{d_{\alpha}/\kappa}h(z,\rho)$ satisfying the initial condition $h(z,0)=t^{\alpha}(z)$. Denote $\xi_{j}=\frac{\partial h}{\partial z^{j}}$, then these equations take the form
@@
where $\Gamma_{ij}^{k}(\rho)=\Gamma_{1/\rho,ij}^{k}$, which are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$. By using the equations (4.11) we can rewrite the equations (4.12) as
@@
Expand $A(\rho)$ and $\zeta{\xi}(\rho)$ into power series of $\rho$ as follows:
$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{0}+\rho\vec{\xi}_{1}+\rho^{2}\vec{\xi}_{2}+\dots,$
@@
$\vec{\xi}_{0}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T},$ (4.14)
then the equation (4.13) is equivalent to
$(A-d_{\alpha})\vec{\xi}_{0}=\mathbf{0},\quad(A-d_{\alpha}-\kappa)\vec{\xi}_{1}=-A_{1}\vec{\xi}_{0},$ (4.15)
$(A-d_{\alpha}-2\kappa)\vec{\xi}_{2}=-A_{2}\vec{\xi}_{0}-A_{1}\vec{\xi}_{1},\dots$ (4.16)
From (3.24) we know that the first equation of (4.15) is satisfied by $\vec{\xi}_{0}$ that is defined by (4.14). Since the eigenvalues of $A$ are positive and less than $\kappa$, we can solve the other linear equations of (4.15) and (4.16) to obtain a unique solution $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ recursively. From the fact that $\vec{\xi}_{0}$ gives a solution of the equations (4.11), and from the compatibility of of the equations of (4.11) and (4.12) we know that $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition. ∎
@@ 4.2. The polynomiality property
The periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\overline{\mathcal{M}}$ given by Proposition 4.1 are linearly independent, so they yield another basis of $\mathcal{S}_{\lambda}$. Thus there exist constants $c_{\alpha}^{\beta}$ and $s^{\beta}$ such that
$\check{x}^{\beta}(z,\rho)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
Since the functions $Z_{r}=Z_{r}(\mathrm{e}^{2\pi i\check{x}^{1}},\dots,\mathrm{e}^{2\pi i\check{x}^{\ell}})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$. Hence we have
@@
It follows from Proposition 4.1 that
$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\rho\to 0.$
Thus we have
@@
$\theta=\inf\{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}~{}\big{|}~{}a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.20)
Suppose $\theta>\theta_{r}$, then it follows from (4.19) that
@@
this is not true due to our definition of proper generators. On the other hand, if $\theta<\theta_{r}$, then from (4.19) we know that
$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\rho\to 0.$ (4.22)
Since $(\theta-\theta_{r})/\kappa<0$, we must have
$\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}=0.$
However, since $t^{1},\cdots,t^{\ell}$ are functionally independent, all those coefficients $a_{n_{1},\cdots,n_{\ell}}$ must vanish. This contradicts with the definition (4.20) of $\theta$. Thus we have $\theta=\theta_{r}$, and it follows that
$z^{r}=\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}.$ (4.23)
In particular, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$. The theorem is proved. ∎
@@
This theorem shows that the proper generators are polynomials of the flat coordinates of $\eta$. It deserves to note that this property also holds true for the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and on the extended affine Weyl groups [12, 14]. For the Frobenius manifold structures constructed in [9] and [12], the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s, while for the ones constructed in [14], $t^{\alpha}$ may involve radical expressions of $z^{j}$’s. For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell})$ for $\ell\geq 5$.
From Theorem 4.2, one can verify that each $g^{\alpha\beta}$ is a rational function of $t^{1},\cdots,t^{\ell}$. Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\overline{\mathcal{M}}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$, here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.53). From the relations (3.38) and (3.53) it follows that we can choose the potential $F$ such that
Therefore, all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
@@
Define the covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$ as (3.59). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\overline{\mathcal{M}}$.
@@
$(t^{1},\cdots,t^{\ell})\colon\overline{\mathcal{M}}\to\mathbb{C}^{\ell}$ (4.25)
is injective. Suppose $p,p^{\prime}\in\overline{\mathcal{M}}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~{}\alpha=1,\cdots,\ell$, then by using Theorem 4.2 we know that
@@
This implies that $\overline{\pi}(p)=\overline{\pi}(p^{\prime})$, where $\overline{\pi}:\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is the canonical projection. Therefore $p\sim p^{\prime}$, i.e., they are identified on $\overline{\mathcal{M}}$. Hence (4.25) is injective, and we proved the corollary. ∎
In conclusion, we have the following commutative diagram:

### v3

@@ 4. Polynomiality Property and the Monodromy Groups
$\mathrm{Stab}_{W}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}$ (4.2)
is a parabolic subgroup of $W(R)$. Moreover, we will construct another generalized Frobenius manifold $\widehat{\mathcal{M}}_{D}$ as a covering space of $\mathcal{M}_{D}$, on which the flat coordinates $t^{1},\dots,t^{\ell}$ are globally defined, and show that its monodromy group is a subgroup of $\mathrm{Mono}(\mathcal{M}_{D})$. To this end, we first need to prove Theorem 4.3 by using properties of the periods of $\mathcal{M}_{D}$, which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
4.1. Periods of the generalized Frobenius manifolds
Let us consider flat coordinates $\tilde{v}(\lambda)$ of the metric $g+\lambda\eta$ of $\mathcal{M}_{D}$. They satisfy the equations
$\nabla_{g+\lambda\eta}\mathrm{d}\tilde{v}(\lambda)=0,$ (4.3)
@@ 4.1. Periods of the generalized Frobenius manifolds
$\mathcal{M}_{\Lambda}=(\mathcal{M}_{D}\times\mathbb{C}^{\times})\setminus{\Sigma_{\Lambda}},$ (4.4)
@@
$\Sigma_{\Lambda}=\{(z,\lambda)\in\mathcal{M}_{D}\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
@@
$\mathscr{Q}(\widetilde{\mathcal{M}}_{\Lambda})=\mathscr{O}(\widetilde{\mathcal{M}}_{\Lambda})/\mathbb{C},$
where $\widetilde{\mathcal{M}}_{\Lambda}$ is the universal covering space of $\mathcal{M}_{\Lambda}$.
In the coordinates $z^{1},\dots,z^{\ell}$, the equation (4.3) reads
$\frac{\partial^{2}\tilde{v}}{\partial z^{i}\partial z^{j}}-\Gamma_{g_{\lambda},ij}^{k}\frac{\partial\tilde{v}}{\partial z^{k}}=0,$ (4.5)
here $\Gamma_{g_{\lambda},ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$. Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a/4\pi^{2}$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$ via the $\varphi_{\lambda}$-transformation. By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ of $\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}$ such that
$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(x^{1}+\theta_{1}c,\dots,x^{\ell}+\theta_{\ell}c\right).$ (4.6)
@@
depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are quasi-homogeneous of degree zero, i.e., they satisfy the equation
$\kappa\lambda\frac{\partial\tilde{v}}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial\tilde{v}}{\partial z^{j}}=0.$ (4.8)
We denote by $\mathcal{S}_{\Lambda}$ the space of periods of $\mathcal{M}_{D}$ that are quasi-homogeneous of degree zero modulo constant functions, then $\mathcal{S}_{\Lambda}$ consists of the solutions to equations (4.5) and (4.8) up to additive constants, hence $\dim_{\mathbb{C}}\mathcal{S}_{\Lambda}=\ell$, and $\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)$ form a basis of $\mathcal{S}_{\Lambda}$.
Now we proceed to find another basis of $\mathcal{S}_{\Lambda}$ associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
Theorem 4.1.
For each quasi-homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, $\alpha=1,\dots,\ell$, there exist a unique quasi-homogeneous period $\tilde{v}\in\mathcal{S}_{\Lambda}$ such that
$\tilde{v}=\tilde{v}^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho),$ (4.9)
where is holomorphic at $\rho=0$ with $h^{\alpha}(z,0)=t^{\alpha}(z)$, or equivalently, $h^{\alpha}(z,\rho)$ has the form
$h^{\alpha}(z,\rho)=t^{\alpha}(z)+\sum_{j=1}^{\infty}h^{\alpha}_{j}(z)\rho^{j}.$ (4.10)
@@
Let us look for a solution of (4.5) and (4.8) of the form $\tilde{v}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ satisfying the initial condition $h^{\alpha}(z,0)=t^{\alpha}(z)$. Denote $\xi_{j}=\frac{\partial h}{\partial z^{j}}$, then these equations take the form
@@
where $\Gamma_{ij}^{k}(\rho)$ are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$. By using the equations (4.11) we can rewrite the equations (4.12) as
@@
Note that $\rho=0$ is a Fuchsian singularity of equation (4.13), so we can solve it by expanding $A(\rho)$ and $\vec{\xi}(\rho)$ into power series of $\rho$ as follows:
$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{(0)}+\rho\vec{\xi}_{(1)}+\rho^{2}\vec{\xi}_{(2)}+\dots,$
@@
$\vec{\xi}_{(0)}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T}.$ (4.14)
Note that $\rho=0$ is a Fuchsian singularity of equation (4.13), its solutions of formal power series always admit a positive radius of convergence. As a result, by expanding equation (4.13) into power series of $\rho$, we only need to solve the linear equations
$(A-d_{\alpha})\vec{\xi}_{(0)}=$ $~\mathbf{0},$ (4.15)
$(A-d_{\alpha}-j\kappa)\vec{\xi}_{(j)}=$ $-\sum_{r=1}^{j}A_{r}\vec{\xi}_{(j-r)},\quad j=1,2,\dots.$ (4.16)
From (3.27) we know that (4.15) automatically holds. Recall from Lemma 3.5 that for each $\beta=1,\dots,\ell$, $\mathrm{Re}(d_{\beta})>0$, so from (3.23) we have $0<\mathrm{Re}(d_{\beta})<\kappa$. This indicates that all the eigenvalues of $A$ have real parts between 0 and $\kappa$, hence we can solve (4.16) to obtain a unique solution $\vec{\xi}_{(1)},\vec{\xi}_{(2)},\dots$ recursively. From the fact that $\vec{\xi}_{(0)}$ gives a solution of the equations (4.11) and the compatibility of the equations of (4.11) and (4.12), we know that $\vec{\xi}_{(1)},\vec{\xi}_{(2)},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition. ∎
@@ 4.2. The polynomiality property
Since periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\mathcal{M}_{D}$ given by Theorem 4.1 are linearly independent (indeed, they are linearly independent near $\rho=0$), they form another basis of $\mathcal{S}_{\Lambda}$. Thus there exist constants $c_{\alpha}^{\beta}$ and $s^{\beta}$ such that
$\check{x}^{\beta}\left(z,1/\rho\right)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
Since the functions $Z_{r}=Z_{r}(\check{x}^{1},\dots,\check{x}^{\ell})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$. Hence we can expand them to get
@@
In order to prove the polynomial dependence of $z^{1},\dots,z^{\ell}$ on $t^{1},\dots,t^{\ell}$, we need the following lemma.
Lemma 4.2.
Suppose $N$ be a positive integer, $b_{1},\dots,b_{N}$ be distinct real numbers, and $p_{1},\dots,p_{N}$ be complex numbers. If
$\lim_{\rho\to 0}\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}=0,$ (4.19)
then $p_{1}=\cdots=p_{N}=0$.
Proof.
Let us define the function $h\colon\mathbb{R}\to\mathbb{C}$ as follows:
$h(s)=\sum_{j=1}^{N}p_{j}e^{-\sqrt{-1}\,b_{j}s}.$
Then from the assumption of the lemma we get
$\lim_{s\to+\infty}h(s)=0,$
from which it follows that
$\lim_{T\to+\infty}\frac{1}{T}\int_{0}^{T}h(s)e^{\sqrt{-1}\,\alpha s}\operatorname{d\!}s=0,\quad\forall\alpha\in\mathbb{R}.$
By taking $\alpha=b_{j}$ for $j=1,\dots,N$, we arrive at $p_{j}=0$. The lemma is proved. ∎
@@
It follows from Theorem 4.1 that
$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\textrm{when}\ \rho\to 0.$
Here and in what follows, the limit $\rho\to 0$ is taken pointwise for each $z\in\mathcal{M}_{D}$. Thus we have
@@
$\theta=\inf\{\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})~\big|~a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.21)
We now show that $\theta=\theta_{r}$. Suppose $\theta>\theta_{r}$, then it follows from (4.20) that
@@
However, this is not true if $z^{r}\not=0$. On the other hand, if $\theta<\theta_{r}$, then from (4.20) we know that
$z^{r}=~\rho^{(\theta-\theta_{r})/\kappa}\sum_{\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})=\theta}\rho^{\sqrt{-1}\,\mathrm{Im}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),$
when $\rho\to 0$. Since each $d_{\alpha}$ admits positive real part, we know that only finitely many vectors $(n_{1},\dots,n_{\ell})\in\mathbb{Z}_{\geq 0}^{\ell}$ satisfy the condition $\mathrm{Re}(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})=\theta$. Suppose $b_{1},\dots,b_{N}$ be all possible distinct values of $\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})$ for those $(n_{1},\dots,n_{\ell})$, then we have
$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}(t^{1},\dots,t^{\ell})+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\textrm{when}\ \rho\to 0,$
where $p_{1},\dots,p_{N}$ are some polynomials of $t^{1},\dots,t^{\ell}$. Since $(\theta-\theta_{r})/\kappa<0$, we must have
$\sum_{j=1}^{N}\rho^{\sqrt{-1}\,b_{j}}p_{j}(t^{1},\dots,t^{\ell})\to 0,\quad\textrm{when}\ \rho\to 0.$ (4.23)
Thus, by using Lemma 4.2 we arrive at $p_{j}=0$ for $j=1,\dots,N$. Since $t^{1},\dots,t^{\ell}$ are functionally independent, all the coefficients $a_{n_{1},\dots,n_{\ell}}$ of the polynomials $p_{j}$ must vanish. This contradicts with the definition (4.21) of $\theta$, so we have $\theta=\theta_{r}$. Now, we can represent $z^{r}$ in the form
$z^{r}$ $=\sum_{\mathrm{Re}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})=\theta_{r}}\rho^{\sqrt{-1}\,\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(1),$ (4.24)
$=p_{0}(t^{1},\dots,t^{\ell})+\sum_{j=1}^{N}\rho^{\sqrt{-1}b_{j}}p_{j}(t^{1},\dots,t^{\ell})+o(1),\quad\textrm{when}\ \rho\to 0.$ (4.25)
Here $b_{1},\dots,b_{N}$ are all distinct nonzero possible values of $\mathrm{Im}(n_{1}d_{1}+\cdots+n_{\ell}d_{\ell})$, and $p_{0},p_{1},\dots,p_{N}$ are polynomials of $t^{1},\dots,t^{\ell}$. By usingLemma 4.2 again, we get $p_{j}=0$ for $j=1,\dots,N$ and $z^{r}=p_{0}$, so $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$. The theorem is proved. ∎
@@
The above theorem shows that the pencil generators are polynomials of the flat coordinates of $\eta$. This property is shared by the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and of the extended affine Weyl groups [12, 14]. The key distinction lies in the inverse relationship: while the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s for the Frobenius manifold structures constructed in [9] and [12], they may involve radical expressions of $z^{j}$’s for the ones constructed in [14]. For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell},\{y_{1},\dots,y_{\ell}\})$ for $\ell\geq 5$.
@@
All degrees $d_{\alpha}$ are positive rational numbers.
Proof.
By Lemma 3.5, it is sufficient to prove that $d_{\alpha}\in\mathbb{Q}$. From Theorem 4.3 we know that each $z^{r}=z^{r}(t),r=1,\dots\ell$ is a quasi-homogeneous polynomial of $t^{1},\dots,t^{\ell}$ of degree $\theta_{r}$, so for any of its monomial term of the form $c(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}$ we have
$n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}.$
Since $z^{1}(t),\dots,z^{\ell}(t)$ are functionally independent, among their monomial terms we can find $\ell$ terms
$f_{k}=c_{k}(t^{1})^{n_{k1}}\cdots(t^{\ell})^{n_{k\ell}},\quad k=1,\dots,\ell$
which are functionally independent, and have degrees $\theta_{m_{1}},\dots,\theta_{m_{\ell}}$. Thus we obtain the following system of linear equations for $d_{1},\dots,d_{\ell}$:
$n_{k1}d_{1}+\cdots+n_{k\ell}d_{\ell}=\theta_{m_{k}},\quad k=1,\dots,\ell.$ (4.26)
Due to the functional independence of $f_{1},\dots,f_{\ell}$, the Jacobian determinant
$\det\begin{pmatrix}\frac{\partial f_{1}}{\partial t^{1}}&\cdots&\frac{\partial f_{1}}{\partial t^{\ell}}\\ \vdots&\ddots&\vdots\\ \frac{\partial f_{\ell}}{\partial t^{1}}&\cdots&\frac{\partial f_{\ell}}{\partial t^{\ell}}\end{pmatrix}=\frac{f_{1}\cdots f_{\ell}}{t^{1}\cdots t^{\ell}}\det\begin{pmatrix}n_{11}&\cdots&n_{1\ell}\\ \vdots&\ddots&\vdots\\ n_{\ell 1}&\cdots&n_{\ell\ell}\end{pmatrix}$
does not vanish identically. It implies that the system of equations (4.26) has non-degenerate coefficient matrix, hence it determines $d_{1},\dots,d_{\ell}$ uniquely in $\mathbb{Q}$. The corollary is proved. ∎
Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widetilde{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$, hence $\pi_{1}(\mathcal{M}_{D})$ acts on those functions. Denote the stable subgroup of this action by
$H=\{\gamma\in\pi_{1}(\mathcal{M}_{D})\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),\ \forall p\in\widetilde{\mathcal{M}}_{D},~\alpha=1,\dots,\ell\},$ (4.27)
then $H$ corresponds to a covering space $\widehat{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$. Concretely, we may define an equivalence relation $\sim$ on $\widetilde{\mathcal{M}}_{D}$ by
Here $\pi\colon\widetilde{\mathcal{M}}_{D}\to\mathcal{M}_{D}$ is the canonical projection. Then we have
$\widehat{\mathcal{M}}_{D}=\widetilde{\mathcal{M}}_{D}/\sim$ (4.29)
with covering map
$\widehat{\pi}:~\widehat{\mathcal{M}}_{D}\to\mathcal{M}_{D}.$
Corollary 4.5.
Define the covering space $\widehat{\mathcal{M}}_{D}$ of $\mathcal{M}_{D}$ as (4.29). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\widehat{\mathcal{M}}_{D}$.
@@
$(t^{1},\cdots,t^{\ell})\colon\widehat{\mathcal{M}}_{D}\to\mathbb{C}^{\ell}$ (4.30)
is injective. Suppose $p,p^{\prime}\in\widehat{\mathcal{M}}_{D}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~\alpha=1,\cdots,\ell$, then by using Theorem 4.3 we know that
@@
This implies that $\widehat{\pi}(p)=\widehat{\pi}(p^{\prime})$, hence $p\sim p^{\prime}$, i.e., they are identified on $\widehat{\mathcal{M}}_{D}$. Therefore, (4.30) is injective. ∎
Corollary 4.6.
Suppose $F$ is the prepotential of $\widehat{\mathcal{M}}_{D}$, then all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
Proof.
From Theorem 4.3, one can verify that each $g^{\alpha\beta}=g(\mathrm{d}t^{\alpha},\mathrm{d}t^{\beta})$ is a rational function of $t^{1},\cdots,t^{\ell}$. Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\widehat{\mathcal{M}}_{D}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$. Here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.56). From the relations (3.41) and (3.56), we can choose the potential $F$ such that
Therefore, all second derivatives of $F$ are rational functions of $t^{1},\dots,t^{\ell}$. ∎
4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and can be extended directly to the case of generalized Frobenius manifolds. Let us first recall this defintion.
Let $M$ be an $\ell$-dimensional generalized Frobenius manifold with flat metric $\eta$ and intersection form $g$. Consider the Gauss-Manin equation
$\nabla_{g}\mathrm{d}v=0.$ (4.32)
for the flat coordinates of $g$. Denote
$\Sigma=\{z\in M\mid\det g(z)=0\},$
then the solutions of (4.32) live on the universal covering $\mathcal{V}$ of $M\setminus\Sigma$. Suppose $v^{1},\dots,v^{\ell}$ be flat coordinates of $g$, then they span, together with constant functions, the solution space of (4.32). For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, $\gamma$ acts on each $v^{\alpha}$ by analytic continuation. Since this action preserves solutions of (4.32), $\gamma^{*}v^{\alpha}$ is still a flat coordinate. Therefore, $\gamma$ yields an affine transformation
$\gamma^{*}v^{\alpha}=A_{\beta}^{\alpha}(\gamma)v^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.33)
on the solution space of (4.32). Here the linear part $A^{\alpha}_{\beta}$ is orthogonal w.r.t. the metric $g$. Let